# DISPATCH - Decentralized Informed Spatial Planning and Assignment of Tasks for Cooperative Heterogeneous Agents

### v2

@@ Abstract
We study fairness in heterogeneous multi-agent systems where tasks vary in preference alignment and urgency. Most existing approaches either assume centralized coordination or largely ignore fairness under partial observability. Distinct from this prior work, we establish a connection between the Eisenberg–Gale (EG) equilibrium convex program and decentralized, partially observable multi-agent learning. Building on this connection, we develop two equilibrium-informed algorithms that integrate fairness and efficiency: (i) a multi-agent reinforcement learning (MARL) framework, EG-MARL, whose training is guided by centralized fair assignment algorithms (EG and a preference-aware Hungarian method); and (ii) a stochastic online optimization mechanism that performs guided exploration and subset-based fair assignment as tasks are discovered.
We evaluate our frameworks across a range of team sizes and assignment formulations against centralized EG, Hungarian, and Min–Max Distance baselines. Both algorithms preserve the fairness–efficiency balance of the Eisenberg–Gale equilibrium under partial observability. EG-MARL achieves near-centralized coordination and reduced travel distances, while the stochastic online mechanism enables real-time allocation with competitive fairness. Together, these results demonstrate that spatially aware EG formulations can effectively guide decentralized coordination in agents with heterogeneous capabilities.
@@ I Introduction
  • Stochastic Online Assignment. We propose a cooperative explore–and–assign mechanism that operates under centralized communication. Whenever k new tasks are discovered, an Eisenberg–Gale (EG) instance is solved over the currently unassigned agents and discovered tasks. Assigned agents immediately begin servicing their targets, while the remaining agents continue exploring, and this process repeats until all tasks are completed. Under this centralized setting, agents solve subset-based Eisenberg–Gale updates as tasks are discovered, providing an upper-bound reference for evaluating EG-MARL under realistic, local communication constraints.
@@ IV EG-MARL Framework
Each agent $i$ observes its own state together with the geometric and semantic attributes of nearby tasks within its sensing range (feature vectors zero-padded for tasks that are not observed).
Let $\mathbf{p}_{i}\in\mathbb{R}^{2}$ and $\mathbf{v}_{i}\in\mathbb{R}^{2}$ denote the agent’s position and velocity in the global coordinate frame, and let $\text{type}_{i}$ and $\text{type}_{j}$ represent the categorical indices of agent $i$ and task $j$, respectively. For every task $j$, the relative position between the task and agent is given by $\mathbf{p}^{j}_{i}=\mathbf{p}^{\text{task}}_{j}-\mathbf{p}_{i}$, and the global preference matrix encodes both the preferences of all agents for each tasks and the individual preference or efficiency $\text{preference}_{ji}$ of agent $i$ toward task $j$. The column vector $\mathbf{p}^{\text{pref}}_{j}\$corresponds to task $j$ ’s preference profile of all agent types, while $w_{j}$ denotes the importance weight associated with the task type. To reflect the dynamic interaction between agents and tasks, we adapt a continuous occupancy variable $\eta_{j}\in[0,1]$ from Aloor et al. [2], defined as $\eta_{j}=1-\min_{i\in\mathcal{N}}\|\mathbf{p}_{i}-\mathbf{p}_{j}^{\text{task}}\|_{2}$, which increases as any agent approaches task $j$. A task history term $h_{j}$ records the last agent to interact with task $j$. For a single task $j$, the observation vector of agent $i$ is defined as:
@@ VI Experiments
We empirically evaluate the proposed EG-MARL and Stochastic Online Assignment frameworks under partial observability across varying team sizes ($N=3,7,10$). Each configuration uses a proportional map size of $2.5\times 2.5$, $2.7\times 2.7$, and $2.9\times 2.9$, respectively, to maintain consistent agent density. The evaluation encompasses three assignment formulations—Eisenberg–Gale, Hungarian, and Min–Max Distance. For the Eisenberg–Gale formulation, we compare decentralized (EG-MARL), online (Stochastic-Online), and centralized implementations, while the Hungarian and Min–Max Distance methods are evaluated only in their centralized forms as baselines.
@@ VI-C Regret under Partial Observability
Assign. Algorithm 3 Agents 7 Agents 10 Agents
EG MARL 0.08 3.99 6.2
Online Assignment 2.22 0.97 3.7
Following Table I, we observe that EG-MARL achieves the lowest regret for smaller teams ($N=3$). As the number of agents increases ($N=7,10$), regret rises for both frameworks, reflecting the greater difficulty of coordinating under partial observability. The Online Assignment performs competitively at larger scales, achieving lower regret than EG-MARL when $N=7$ and $N=10$. Overall, EG-MARL maintains more consistent proximity to the centralized optimum, while the online method demonstrates better scalability.
@@ VI-D Fairness and Efficiency Analysis
Table II and Figure 3 summarize results across all configurations. For the Eisenberg–Gale (EG) formulation, we compare EG-MARL, Online Assignment, and the Centralized Oracle. The Hungarian and Min–Max Distance methods serve as centralized baselines without corresponding MARL or online counterparts. Metrics include completion time ($T$), total distance ($D$), and fairness ($F(\rho)$), where lower $T$ and $D$ and higher $F(\rho)$ indicate better performance.
@@
Figure 3: Fairness–Efficiency tradeoff across all algorithms, shown using $F(\rho)$ versus completion time for 3, 7, and 10 agents.
@@
For the 7 and 10-agent cases, the EG optimization reveals complementary strengths across the three methods: EG-MARL achieves the fastest completion times, the Stochastic-Online approach maintains the highest fairness, and the Centralized oracle yields the lowest travel distances. This diversity of strengths highlights the trade-offs among scalability, fairness, and efficiency that emerge under different coordination paradigms.
@@
Regret is lowest for moderate subsets (k=3k=355), which balance exploration time with the assignment optimality gap from the centralized solution. As $k$ increases, completion time $T$ decreases monotonically since larger subsets more closely approximate the centralized assignment; when $k=7$, the allocation becomes identical to the centralized EG solution, minimizing idle time but increasing travel distance. In both fairness measures, $F(\rho)$ and Jain’s Index $J(\rho)$, remain largely stable across all subset sizes $k$, indicating that the online assignment preserves equilibrium balance regardless of how many agents are jointly considered. Minor variations at higher $k$ reflect small sensitivity but do not alter fairness trend.
@@ VI-E Ablation Study
MPE Distance Penalty In the ablation study, we remove the EG fairness-shaping reward that originally guided agents toward their equilibrium-assigned tasks. Instead, the ablated reward uses a standard MPE-style distance-shaping term that is for each task $j$, the environment computes the distances from all agents to that task and subtracts the minimum distance:
@@
We also evaluate the MPE baseline reported without the workload–progress reward. In addition, we construct a second EG-MARL variant that removes the exploration reward for the ablation study. Table IV summarizes the 7-agent ablation results.
TABLE IV: Ablation results for the 7-agent setting: EG-MARL vs. Ablated (MPE reward only).
@@
We observe that removing the EG fairness-shaping increases both completion time and travel distance. Despite relying solely on distance-based shaping, the ”MPE Baseline” maintains reasonable fairness, though it consistently underperforms the EG-MARL model. This indicates that the equilibrium-guided EG reward accelerates convergence while strengthening fairness.
We hypothesize that although the “No Exploration Reward” variant retains a strong navigation signal, the absence of exploration incentives leads agents to prematurely commit to early-discovered tasks. While its completion time and total distance remain comparable to those of full EG-MARL, this similarity may partly reflect statistical variability in the experimental setting rather than a true behavioral equivalence.
@@ VII Assumptions and Limitations
Our formulation makes several modeling assumptions. We consider one-to-one assignments with no preemption, and tasks are static in our experiments. EG-MARL assumes local, range-limited communication with multi-hop propagation, while the Online Assignment method operates under full centralized communication and serves as a baseline. The online subset enumeration is tractable only for small teams, and heuristic approximations would be required for larger deployments.

### v3

@@ Abstract
We study fairness in heterogeneous multi-agent systems where tasks vary in preference alignment and urgency. Most existing approaches either assume centralized coordination or largely ignore fairness under partial observability. Distinct from this prior work, we establish a connection between the Eisenberg–Gale (EG) equilibrium convex program and decentralized, partially observable multi-agent learning. Building on this connection, we develop two equilibrium-informed algorithms that integrate fairness and efficiency: (i) a multi-agent reinforcement learning (MARL) framework, EG-MARL, whose training is guided by centralized EG equilibrium assignment algorithm; and (ii) a stochastic online optimization mechanism that performs guided exploration and subset-based fair assignment as tasks are discovered.
We evaluate on Multi-Agent Particle Environment (MPE) simulations across varying team sizes against centralized EG, Hungarian, and Min–Max distance baselines, and also present a Webots-based warehouse proof-of-concept with heterogeneous robots. Both methods preserve the fairness–efficiency balance of the EG solution under partial observability, with EG-MARL achieving near-centralized coordination and reduced travel distances, and the online mechanism enabling real-time allocation with competitive fairness.
@@ I Introduction
  • Stochastic Online Assignment. We propose a cooperative explore–and–assign mechanism that operates under centralized communication. Whenever k new tasks are discovered, an Eisenberg–Gale (EG) instance is solved over the currently unassigned agents and discovered tasks. Assigned agents immediately begin servicing their targets, while the remaining agents continue exploring, and this process repeats until all tasks are completed.
We validate our approach in Multi-Agent Particle Environment (MPE) simulations across multiple team sizes and in a Webots warehouse proof-of-concept with heterogeneous robots.
@@ IV EG-MARL Framework
Each agent $i$ observes its state with the geometric and semantic attributes of nearby tasks within its sensing range (feature vectors zero-padded for tasks that are not observed).
A task becomes observable only via local sensing or communication.
Let $\mathbf{p}_{i}\in\mathbb{R}^{2}$ and $\mathbf{v}_{i}\in\mathbb{R}^{2}$ denote the agent’s position and velocity in the global coordinate frame, and let $\text{type}_{i}$ and $\text{type}_{j}$ represent the categorical indices of agent $i$ and task $j$, respectively. For every task $j$, the relative position between the task and agent is given by $\mathbf{p}^{j}_{i}=\mathbf{p}^{\text{task}}_{j}-\mathbf{p}_{i}$, and the global preference matrix encodes the preference profile of all tasks all agents and the preference or efficiency $\text{preference}_{ji}$ of agent $i$ toward task $j$. The column vector $\mathbf{p}^{\text{pref}}_{j}\$corresponds to task $j$ ’s preference profile of all agent types, while $w_{j}$ denotes the importance weight associated with the task type. To reflect the dynamic interaction between agents and tasks, we adapt a continuous occupancy variable $\eta_{j}\in[0,1]$ from Aloor et al. [2], defined as $\eta_{j}=1-\min_{i\in\mathcal{N}}\|\mathbf{p}_{i}-\mathbf{p}_{j}^{\text{task}}\|_{2}$, which increases as any agent approaches task $j$. A task history term $h_{j}$ records the last agent to interact with task $j$. For a single task $j$, the observation vector of agent $i$ is defined as:
@@ V Stochastic Online Assignment
The primary role of the stochastic online EG mechanism is to serve as a strong online baseline with centralized communication against which to evaluate EG-MARL, rather than as a directly deployable algorithm at very large scales. The subset enumeration step is combinatorial in the team size, but the resulting policies are highly informative: they reveal how close decentralized learned policies come to an online EG-guided allocation with full centralized coordination. In larger systems, the same framework could be combined with heuristic or sampling-based subset selection strategies (e.g., random sampling of candidate subsets or prioritizing subsets based on recent travel costs), which we leave for future work.
@@ VI Experiments
We empirically evaluate the proposed EG-MARL and Stochastic Online Assignment frameworks under partial observability across varying team sizes ($N=3,7,10,15$). Each configuration uses a proportional map size of $2.5\times 2.5$, $2.7\times 2.7$, $2.9\times 2.9$, and $3.2\times 3.2$ respectively, to maintain consistent agent density. The evaluation encompasses three assignment formulations—Eisenberg–Gale, Hungarian, and Min–Max Distance. For the Eisenberg–Gale formulation, we compare decentralized (EG-MARL), online (Stochastic-Online), and centralized implementations, while the Hungarian and Min–Max Distance methods are evaluated only in their centralized forms as baselines.
@@ VI-C Regret under Partial Observability
Assign. Algorithm 3 Agents 7 Agents 10 Agents 15 Agents
EG MARL 0.08 3.99 6.2 9.88
Online 2.22 0.97 3.7 7.05
Following Table I, we observe that EG-MARL achieves the lowest regret for smaller teams ($N=3$). As the number of agents increases ($N=7,10,15$), regret rises for both frameworks, reflecting the greater difficulty of coordinating under partial observability. The Online Assignment performs competitively at $N=7$, $N=10$, and $N=15$.
@@ VI-D Fairness and Efficiency Analysis
Table II and Figure 3 summarize results across all configurations. For the Eisenberg–Gale (EG) formulation, we compare EG-MARL, Online Assignment, and the Centralized Oracle. The Hungarian and Min–Max Distance methods serve as centralized baselines. Metrics include completion time ($T$), total distance ($D$), and fairness ($F(\rho)$), where lower $T$ and $D$ and higher $F(\rho)$ indicate better performance.
@@
15 EG MARL 3.01 26.19 1.64 0.717
Online (k=6) 6.90 46.82 1.63 0.715
Centralized 3.10 25.25 1.68 0.727
Hungarian Centralized 6.14 24.9 1.195 0.572
Min–Max Centralized 6.77 25.4 1.196 0.568
Figure 3: Fairness–Efficiency tradeoff across all algorithms for 3, 7, 10, and 15 agents.
@@
For the 7, 10, and 15-agent cases, the EG optimization reveals complementary strengths across the three methods: EG-MARL achieves the fastest completion times, the Stochastic-Online approach maintains the highest fairness, and the Centralized oracle yields the lowest travel distances. These strengths highlight the trade-offs among scalability, fairness, and efficiency that emerge under different coordination paradigms.
@@
Regret is lowest for moderate subsets (k=3k=355), which balance exploration time with the assignment optimality gap from the centralized solution. As $k$ increases, completion time $T$ decreases monotonically since larger subsets more closely approximate the centralized assignment; when $k=7$, the allocation becomes identical to the centralized EG solution, minimizing idle time but increasing travel distance. In both fairness measures, $F(\rho)$ and Jain’s Index $J(\rho)$, remain largely stable across all subset sizes $k$, indicating that the online assignment preserves equilibrium balance regardless of how many agents are jointly considered. Minor variations at higher $k$ reflect small sensitivity but do not alter the fairness trend.
@@ VI-E Ablation Study
Simple Spread MPE
In the ablation study, we remove the EG fairness-shaping reward that originally guided agents toward their equilibrium-assigned tasks. Instead, the ablated reward uses a standard MPE-style distance-shaping term that, for each task $j$, computes the minimum distance from any agent to that task: Intuitively, this mirrors the MPE ’spread’ reward: using the minimum agent–task distance encourages the team to spread out so that at least one agent is close to every task. However, because this shaping ignores task priorities, workloads, and heterogeneous preferences, it promotes spatial coverage rather than fairness‑aware assignment.
@@
We also evaluate this baseline without the workload–progress reward. While all other rewards remain the same as EG-MARL.
No Exploration Reward
We construct a second EG-MARL variant that removes the exploration reward.
TABLE IV: Ablation results for the 7-agent setting: EG-MARL vs. Simple Spread MPE. vs. No Exploration Reward.
@@
We observe from Table IV that removing the EG fairness-shaping increases both completion time and travel distance. Despite relying solely on distance-based shaping, the Simple Spread MPE Baseline maintains reasonable fairness, though it consistently underperforms the EG-MARL model.
While the No Exploration Reward variant yielded completion time and total distance remaining comparable to those of the full EG-MARL model, its fairness metrics degraded slightly. We hypothesize that, without an exploration incentive, agents prematurely commit to early-discovered tasks, which in turn reduces fairness.
VI-F Webots Validation in a Warehouse Environment
Experimental Setup
We use the centralized assignemnt method to demonstrate a use-case of our framework in a Webots warehouse-like scene containing a cabinet, pallets, box and gas container. Five heterogeneous differential-drive robots serve as agents, and a stationary supervisor centrally computes assignments to fice differenst tasks. Each robot and workstation has an associated type that determines skill preference and workload priority. The supervisor collects robot positions, workstation locations, and obstacles, and constructs utilities using distance-discounted preference scores.
The Webots supervisor first constructs a two-dimensional occupancy grid of the warehouse and computes the shortest collision-free path from each robot $i$ to each workstation $j$ using a grid-based A* planner. These path lengths define the distance-aware utilities used in the Eisenberg–Gale assignment, which determines the robot–workstation matching. During execution, the supervisor streams the next waypoint of each robot’s precomputed path through low-bandwidth emitter messages, while the robots rely on their onboard go-to-goal controllers to follow the path and avoid obstacles.
Case Study: Heterogeneous Robots and Tasks
To model heterogeneous robot capabilities and the varying urgency of warehouse tasks, we define skill–task compatibility as follows:
  • Summit-XL Steel: A heavy-duty platform with the strongest capability for pallet and other high-load tasks.
  • youBot: Well-suited for delicate manipulation tasks such as handling gas canisters due to its dexterous arm.
  • TurtleBot3 Burger: A lightweight, agile platform designed for navigating narrow spaces and interacting with drawers, cabinets, and enclosed storage.
  • Pioneer 3-AT: A robust all-terrain robot with sreliable performance across moderately heavy warehouse tasks.
  • Pioneer 3-DX: Optimized for fast navigation and high-throughput roles, best with conveyor and open-flow stations.
Task-weight (budget) values further encode urgency—high-throughput or safety-critical stations, such as the conveyor or gas canister storage, receive larger weights, while low‑impact stations, such as cardboard boxes, receive lower ones.
Figure 4 illustrates the resulting assignment in our Webots deployment. The Summit-XL Steel is assigned to the wooden pallet stack, the TurtleBot3 Burger to the gas canister, the Pioneer 3-DX to the conveyor platform, the Pioneer 3-AT to the cabinet, and the youBot to the cardboard box.
Figure 4: Eisenberg–Gale assignment in Webots. The Summit-XL Steel is assigned to the pallet stack, the TurtleBot3 Burger to the cabinet, the Pioneer 3-DX to the conveyor, the youBot to the gas canister, and the Pioneer 3-AT to the cardboard box.
Performance Evaluation
We evaluate the Webots warehouse deployment using the same fairness and efficiency metrics as in our MPE experiments.
TABLE V: Webots Evaluation Metrics for EG, Hungarian, and Min–Max. Bold values indicate the best performance per column.
Method $T$ ($\downarrow$) $D$ ($\downarrow$) $F(\rho)$ ($\uparrow$) $J(\rho)$ ($\uparrow$)
EG 54.53 36.82 1.710 0.745
Hungarian 51.78 34.27 1.259 0.613
Min–Max 70.85 36.86 1.438 0.674
Summary of Findings
In this case study, the centralized Eisenberg–Gale (EG) allocation achieves the strongest overall balance between fairness and efficiency. EG attains the highest fairness scores ($F(\rho)$ and $J(\rho)$) while also producing the competitive completion time among the three assignment methods. The Hungarian method, despite achieving the best completion time and competitive travel distance, exhibits substantially lower fairness due to prioritizing preference alignment without regard to the priority weight across tasks. The Min–Max baseline yields the shortest total distance traveled but shows lower fairness compared to EG, as minimizing worst-case travel does not incorporate task urgency or skill–task compatibility. These results are consistent with our simulation findings and demonstrate that EG equilibrium-guided assignment provides more equitable task outcomes while maintaining competitive spatial efficiency in realistic heterogeneous multi-robot deployments. This warehouse scenario mirrors our navigation-and-service formulation: tasks correspond to shelf locations with item requests, and heterogeneous robots must navigate around obstacles to service them. Although this experiment uses centralized supervision rather than learned decentralized policies, it demonstrates that the fairness–efficiency trade-offs observed in MPE simulations carry over to a realistic, heterogeneous multi-robot warehouse setting.
@@ VII Assumptions and Limitations
Our formulation makes several modeling assumptions. We consider one-to-one assignments with no preemption, and tasks are static in our experiments. EG-MARL assumes local, range-limited communication with multi-hop propagation, while the Online Assignment method operates under full centralized communication and serves as a baseline. We remark that the computational bottleneck of the Stochastic Online Assignment is dominated by the combinatorial enumeration term $\binom{N}{k}$, resulting in a complexity of $T=\mathcal{O}\left(\binom{N}{k}\cdot\text{Poly}(k)\right)$. This exponential dependence on the team size $N$ limits its practical use to systems with a small number of agents. In addition, we recognize the statistical bias from the testing samples. Finally, we acknowledge that the test environments for different team sizes were not held constant, which may introduce sample bias in the results across settings.
Finally, we clarify that EG-MARL is not an approximation algorithm for the underlying Dec-POMDP, but rather an empirically EG-guided learning framework that uses the equilibrium solution only as a supervisory signal during training.